### v2

# DEM: Distribution Edited Model for Training with Mixed Data Distributions

Abstract
Training with mixed data distributions is a common and important part of creating multi-task and instruction-following models. The diversity of the data distributions and cost of joint training makes the optimization procedure extremely challenging. Data mixing methods partially address this problem, albeit having a sub-optimal performance across data sources and require multiple expensive training runs. In this paper, we propose a simple and efficient alternative for better optimization of the data sources by combining models individually trained on each data source with the base model using basic element-wise vector operations. The resulting model, namely Distribution Edited Model (DEM), is $\times1$ cheaper than standard data mixing and outperforms strong baselines on a variety of benchmarks, yielding upto 6.2% improvement on MMLU, 11.5% on BBH, 16.1% on DROP, 6% on MathQA, and 9.3% on HELM with models of size 3B to 13B. Notably, DEM does not require full re-training when modifying a single data-source, thus making it very flexible and scalable for training with diverse data sources.
1 Introduction
Figure 1: The Distribution Edited Model ($\Theta_{D}$) results from fine-tuning a pretrained model ($\Theta$) on $n$ individual data distributions ($D_{i}$) and combining the resulting models with basic element-wise vector operations. Here, the combination is achieved by extracting distribution vectors ($\Delta\Theta_{D_{i}}$), multiplying them by weight coefficients ($\omega_{i}$), and adding their weighted sum to the base model.
Large Language Models (LLM) go through an extensive pretraining on billions or trillions of tokens (Brown et al., 2020; Zhang et al., 2022; Raffel et al., 2020; Touvron et al., 2023a, b; Geng and Liu, 2023), but they typically require supervised fine-tuning on diverse instruction-following datasets for properly following human instructions Ouyang et al. (2022); Sanh et al. (2022); Iyer et al. (2022); Chung et al. (2024). Supervised training is crucial for ensuring that generated outputs meet user expectations and perform well on downstream tasks Radford et al. (2019); Gururangan et al. (2020).
The datasets for supervised training are often of different sizes and follow different distributions. Recent state-of-the-art fine-tuning approaches Iyer et al. (2022); Chung et al. (2024) demonstrate that training on multiple data distributions requires careful tuning of the mixing weights for each data source to capture the combined distribution and improve downstream task performance. Tuning these weights in a data-mixing approach is a computationally expensive process. Although, there are techniques to speed-up the search (Xie et al., 2023; Albalak et al., 2023), the process remains time-consuming. Moreover, when one or more new datasets are introduced, the weights for each dataset need to be re-tuned. This requirement makes the data-mixing approach inflexible and hard to maintain in a production environment.
To address these challenges when fine-tuning an LLM on a set of diverse data distributions, we propose a simple and efficient approach that combines individually trained versions of the base model using element-wise vector operations. Our method focuses on the challenging setting of combining diverse data distributions that correspond to multiple tasks from different domains such as math, reasoning, conversations and coding. In particular, our goal is to better capture a diverse data distributions as opposed to editing the model on a single downstream task Ilharco et al. (2022a); Schumann et al. (2024). Hence, we call resulting model Distribution Edited Model (DEM, shown in Figure 1). Our experiments on a variety of downstream tasks show that DEM is an effective, highly capable and low cost alternative to the models trained using data mixing methods.
The primary benefit of the proposed approach is its ability to efficiently identify the optimal combination of data sources for training a model. Instead of exhaustively training and validating on all possible combinations of data sources, which can be computationally expensive, we take a more streamlined approach. First, we finetune the original model on each individual data source independently with early stopping to obtain the optimal model. Second, we extract source distribution vectors by subtracting the original model from the finetuned ones. Lastly, we create the final model by adding a weighted combination of these distribution vectors to the base model, allowing it to capture the joint distribution of different data sources in a controlled manner while enabling incremental updates with new datasets.
Our contributions can be summarized as follows:
  • We propose a simple and efficient approach for training models on diverse data distributions that offers a flexible way for tuning the contributions of each data source individually without the need of full data re-training (Section 4).
  • We show that DEM reduces the training cost by $\times1$ while improving the model performance. Compared to standard data mixing approaches, DEM yields up to 6.2% improvement on MMLU, 11.5% on BBH, 16.1% on DROP, 6% on MathQA and 9.3% on HELM with 3B, 7B, and 13B models.
  • We perform an exhaustive analysis of the properties of the distribution vectors and their corresponding models, finding that DEM is better aligned with the individual models than baseline while remaining close to the original model.
2 Related Work
Multi-task Fine Tuning
Instruction-based multi-task fine-tuning of language models has been previously shown to improve both zero and few-shot performance on unseen tasks  (Wei et al., 2022a; Sanh et al., 2022). Instruction-tuning data can be sourced from diverse task categories (such as math, reasoning, dialog etc), and the model performance is often sensitive to the data-mixing strategy. For example, both  (Chung et al., 2024) and  (Iyer et al., 2022) carefully tune the data-mixing weights for various training data sources.
Hyperparameter tuning of data-mixing weights is a compute intensive process, and methods such as DoReMi  (Xie et al., 2023) and Online Data Mixing  (Albalak et al., 2023) have been proposed to speed-up the process for pretrained data-mixing either through a proxy-model training or through a multi-armed bandit approach respectively. Renduchintala et al. (2024) used a submodular function to assign importance scores to tasks which are then used to determine the mixture weights. Li et al. (2024) built a framework to find multiple diverse solutions in the Pareto front of many objectives. In this work, we propose an alternative strategy for training with multiple data sources by using vector arithmetic to combine models fine-tuned on individual datasets, rather than mixing training data in specific proportions.
Model Weight Interpolation
Recently, model weight interpolation and task arithmetic techniques have been shown to improve the performance of pre-trained models on: single-task Izmailov et al. (2018); Matena and Raffel (2022); Yüce et al. (2022); Wortsman et al. (2022b) and multi-task Ilharco et al. (2022b, a); Li et al. (2022); Wortsman et al. (2022a); Yadav et al. (2023); Daheim et al. (2024), out-of-domain generalization Arpit et al. (2022); Rame et al. (2022); Jin et al. (2023); Ramé et al. (2023); Cha et al. (2024), and federated learning McMahan et al. (2017); Li et al. (2020).
Going beyond simple weight averaging, (Matena and Raffel, 2022) explored merging using Fisher-weighted averaging for improving single-task model performance by leveraging other auxiliary tasks. Ilharco et al. (2022a) presented a model merging technique based on producing task vectors and performing arithmetic operations, such as addition, subtraction to obtain a multitask checkpoint and ‘forget’ unwanted behavior. Daheim et al. (2024) proposed a new uncertainty-based correction of the task vector coefficients to improve the performance by reducing the model mismatch.
While previous work focused on classification tasks in NLP or vision, we extend vector-arithmetic-based model editing to multi-task fine-tuning on diverse data distributions. Our results show that the proposed approach outperforms and is more efficient than data-mixed fine-tuning.
3 Background: Data mixing
Let us consider a pretrained language model with parameters $\Theta$, and $D_{1},D_{2},...,D_{n}$ denote $n$ different supervised fine tuning datasets. Each dataset can consist of a single or multiple tasks. The exact tasks may have an overlap between these datasets, however, the corresponding samples are unique to each dataset. Standard data mixing Chung et al. (2024); Iyer et al. (2022) methods create training batches by performing a weighted sampling from each of the training datasets $D_{i}$. The goal is to learn a joint data distribution that can span all training datasets.
4 Proposed Approach: Distribution Edited Model (DEM)
In contrast to standard data mixing, we propose to learn each data distribution separately and combine them post training. In the following subsections, we present two variants of that lead to a Distribution Edited Model that achieves this goal.
4.1 Combined Distribution Vectors
Let us assume a set of training data sources ($D_{1},D_{2},...,D_{n}$). First, we fine tune our pretrained model ($\Theta$) on each of these $n$ datasets separately, with a different set of hyper-parameters (chosen for optimal validation loss). The corresponding fine-tuned models are noted as $\Theta_{D_{1}},\Theta_{D_{2}},...,\Theta_{D_{n}}$. Next, we define a data distribution vector (DV) $\Delta\Theta_{D_{i}}$ (corresponding to the dataset $D_{i}$) as the element-wise difference of parameters between the pretrained model $\Theta$ and a fine-tuned model $\Theta_{D_{i}}$, following a similar approach as presented in Ilharco et al. (2022a).
$\leavevmode\nobreak\ \Delta\Theta_{D_{i}}=\Theta_{D_{i}}-\Theta,$ (1)
Instead of task specific model editing, as in prior work, we focus on a mixture of large number of diverse NLP downstream tasks. These different tasks are represented with their own data distribution and we investigate how to combine different data DVs that we can extract by fine tuning the pretrained model using data from different distributions.
Lastly, we obtain a mixed data DV by computing a weighted combination of each $\Delta\Theta_{D_{i}}$ with corresponding weights $\omega_{i}\in\mathbb{R}$. Finally we add the pretrained model $\Theta$ to obtain our Distribution Edited Model (DEM) as follows:
$\leavevmode\nobreak\ \Theta_{D}=\Theta+\sum_{i=1}^{n}\omega_{i}\Delta\Theta_{D_{i}}.$ (2)
4.2 Model Interpolation
Another way to combine the finetuned models ($\Theta_{D_{i}}$) is through model weight interpolation. In this case, we do not extract data distribution vectors ($\Delta\Theta_{D_{i}}$), but rather use the finetuned models directly that capture information about the data distribution. Specifically, we take a weighted average of all the fine tuned models ($\Theta_{D_{i}}$) where the weights, $\omega_{i}\in\mathbb{R}$ sum to 1. More formally,
$\leavevmode\nobreak\ \Theta_{D}=\sum_{i=1}^{n}\omega_{i}\Theta_{D_{i}};\quad s.t.\quad\sum_{i=1}^{n}\omega_{i}=1.$ (3)
Note that, Eq 3 is a special case of Eq 2 when $\sum_{i=1}^{n}\omega_{i}=1$. DEM using distribution vector (Eq 2) provides more flexibility in terms of choosing $\omega_{i}$ per data source which can yield further performance improvement (Section 6.1).
4.3 Computational Cost
To better understand the advantages of DEM over the data mixing approach we derive a formula to measure the cost for each method. Let us assume we have $n$ different data sources and $m$ number of weights per data source. The hyperparameter search space for both the approaches will have a total of $m^{n}$ weight combinations. Intuitively, searching for data mixing weights is comparatively more expensive than DEM since full data re-training is required for validating each weight combination. On the other hand, DEM requires finding the optimal weights after individual training using only validation for each weight combination.
To formalize this, assume $T$ and $V$ as the average number training and validation steps respectively. The computational complexity for the weighted data mixing will be $O(m^{n}(T+V))$ and for the proposed DEM approach will be $O(n(T+V)+(m^{n}V))$. We can clearly see that $O(m^{n}(T+V))\geq O(n(T+V)+(m^{n}V))$, and with DEM we reduce the number of training run by a factor of $m^{n}/n$.
Additionally, we can compare the exact training and validation cost of our proposed DEM approach with the baseline. Assuming $k$ steps of training or validation and each step takes $t$ seconds, we can define the cost ($c$) in gpu-hours as follows:
$c=(k*t*g)/3600$ (4)
where $g$ is the total number of GPUs used. The exact cost for training ($c_{train}$) and validation ($c_{val}$) depends on the corresponding values of $k$, $t$ and $g$ and generally $c_{train}\gg c_{val}$.
5 Experimental Setup
5.1 Dataset
Here, we list the fine-tuning datasets, we use to enhance instruction following capability of our base pre-trained LLM. Previous work has shown that they improve the instruction following capabilities of the models (Chung et al., 2024; Iyer et al., 2022; Gupta et al., 2022; Amini et al., 2019; Sanh et al., 2022).
  • Chain of Thoughts (CoT) Wei et al. (2022b): The CoT mixture Chung et al. (2024) consists of nine datasets with manually written CoT annotations. Each task in these nine datasets have ten manually composed instruction templates, and the span arithmetic reasoning, multi-hop reasoning, and natural language inference.
  • Math QA Amini et al. (2019): This dataset consists of 37K math-based multiple-choice word problems. The problem set includes geometry, counting, probability, physics, gain-loss and other general math topics.
  • Public Pool of Prompts (P3) Sanh et al. (2022): P3 (Public Pool of Prompts) is a collection of prompted English datasets for a diverse set of NLP tasks, where each sample consists of a prompted input and a target text. Prompts can be considered as a functions that map an example from a dataset to a natural language input and target output. Promptsource Bach et al. (2022) is used to interactively create prompts and gather prompt-specific metadata like evaluation metrics. As of writing of this paper, over 2,000 prompts from 270+ datasets are publicly available on Promptsource.
  • Instruct Dial (InstDial) Gupta et al. (2022): This is an instruction tuning dataset designed for dialogues. It consists of 48 different dialogue tasks from 59 open dialogue datasets which is unified in text-to-text format suitable for decoder LLMs. It has been shown improve model performance in unseen datasets, specially for dialogue related tasks.
  • Super Natural Instructions (SNI) Wang et al. (2022): This dataset consists of 1,616 diverse NLP tasks in text-to-text format with instructions written by experts. It covers 76 distinct task types, including but not limited to text composition, infilling, extraction, classification, sequence tagging and paraphrasing.
5.2 Model Architecture
We use OpenLLaMA Geng and Liu (2023) as our base LLM, which is trained on 1T tokens from the RedPajama Dataset Computer (2023). It follows the same architecture as the LLaMA model Touvron et al. (2023a) – a decoder-only LLM with rotary positional embedding, SwiGLU activations and RMS Norm for pre-normalization. In our experiments, we cover three different sized models: 3B, 7B and 13B (see Table 1). We carry all ablations with the 7B model, while the 3B and 13B models are used to show generalization of the proposed approach to other sizes. The experimental results show that the properties of DEM are present across different model sizes.
# Params Context Dims # Heads # Layers
3B 2048 3200 32 26
7B 2048 4096 32 40
13B 2048 5120 40 40
Table 1: Characteristics of different OpenLLaMA model sizes used in our experiments.
5.3 Training
We fine-tune the OpenLLaMA model on all instruction following datasets (Section 5.1), both separately and jointly. We use AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay of 0.05, gradient clipping of 1 and a constant learning rate of 2e-5 with a 2000 step warmup. We also adjust batch size for different datasets based on the validation loss (see Appendix A for details). We use a greedy sample packing approach to fit multiple training samples into a single batch sample efficiently, padding to the max sequence length without overflowing into the next sample of a batch. To select the optimal mixing weights for DEM (Eq 2), we perform a grid search over $\omega_{i}$ values. For each coefficient combination we evaluate the validation losses on the five datasets (Section 5.1), and select the model that minimizes their average (see Section C for details). We use an equal weight of $\omega_{i}=0.25$ for all datasets in our experiments.
5.4 Evaluation Framework
We evaluate the instruction following capability of the models using three publicly available benchmarks, namely InstructEval (Chia et al., 2024), LM-evaluation-harness (Gao et al., 2024) and HELM (Liang et al., 2023). To have a holistic evaluation, we choose a diverse set of held-out datasets: (i) from InstructEval – MMLU, Big-Bench Hard and DROP, (ii) from LM-evaluation-harness – MathQA, and (iii) from HELM – twenty sets from six diverse task-groups – Classification, ClosedbookQA, OpenbookQA, Math, Reasoning and Conversational (see Table 3). We perform 5-shot evaluation on MMLU and HELM, and 3-shot evaluation on BBH and DROP, inline with the standardized setup and previous work.
5.5 Baseline Models
The pre-trained OpenLLaMA serves as the non instruction-tuned baseline for evaluation. Our primary instruction-tuned baseline is data mixing – the model is fine-tuned using a weighted mixture of 5 diverse datasets as described in 5.1 following Chung et al. (2024); Iyer et al. (2022) which has been shown to produce SOTA performance with large scale diverse datasets. This model requires finding the optimal weights corresponding to each training dataset such that the validation loss reaches optimal value for each dataset at similar number of training steps. We experimented with several combinations of weights and chose the one that leads to the smallest validation loss (see Appendix B for details). Additionally, we create a simpler baseline where we concatenate all 5 training datasets and the samples are shuffled randomly during training. This technique is more cost-effective than the standard data mixing approach, as it does not require any weight optimization.
6 Experimental Analysis
6.1 Downstream Task Performance
Models MMLU BBH DROP MathQA
Open LLaMA 40.31 32.84 24.38 27.71
LLaMA Touvron et al. (2023a) 35.10 30.30 - -
LLaMA2 Touvron et al. (2023b) 45.30 32.60 - -
FlanPaLM (8B) Chung et al. (2024) 49.3 36.4 - -
OPT-IML (30B) Iyer et al. (2022) 43.2 30.9 - -
OPT-IML (175B) Iyer et al. (2022) 47.1 35.7 - -
CoT 41.67 33.98 24.20 29.31
Math QA 39.71 32.70 24.31 25.03
P3 35.69 14.00 23.29 27.14
InstDial 39.31 23.09 21.81 26.60
SNI 46.55 35.88 34.53 28.31
Data Mixing 47.77 36.38 32.71 30.35
Concatenated Datasets 43.43 21.34 23.21 27.71
DEM - Interpolation (Ours) 50.14 40.11 36.31 31.22
DEM - Distribution Vector (Ours) 50.74 40.56 37.96 32.16
Table 2: Downstream task performance of models trained on different instruction following datasets (Section 5.1). We compare it with different pretrained and fine-tuned baselines (Section 5.5) and our proposed approach in Section 4. The models are of size 7B, unless specified. The performance numbers for models with citation are taken from the corresponding paper, rest are evaluated using InstructEval and LM-evaluation-harness.
In this section, we first use the Instruct-Eval framework to evaluate the performance of both the pre-trained and fine-tuned models. The performance on MMLU, BBH and DROP is shown in Table 2.111The low performance on BBH after training on P3 is inline with previous findings Iyer et al. (2022). The T0pp 11B’s Sanh et al. (2022) accuracy is 13.0, after being fine-tuned only on the P3 dataset. In addition to the pretrained OpenLLaMA model, we show the performance of LLaMA Touvron et al. (2023a) and LLaMA2 Touvron et al. (2023b) of same size as a reference. We also include three other supervised fine tuned models of larger sizes, namely FlanPaLM (8B) and OPT-IML (30, 175B).
We present the performance of fine tuned models on each dataset separately and observe that the performance degrades compared to Open-LLaMA model for P3, InstDial and MathQA. On the other hand, we observe significant improvement with CoT and SNI datasets in all four task families. We compare these models with data mixing baseline and note that it performs significantly better than the pretrained OpenLLaMA, while the improvement compared to the best single dataset fine-tuned model (i.e. SNI) is much smaller, even worse for DROP. The concatenated datasets baseline performs significantly worse than data mixing method, only improving for MMLU compared to OpenLLaMA and significantly worse than SNI fine-tuned model. This highlights the importance of choosing the optimal weights for data mixing and training a strong baseline.
Models Classification Closedbook QA OpenBook QA Math Reasoning Conversational
OpenLLaMA 49.68 23.21 48.55 10.45 50.40 33.03
Data Mixing 56.52 28.71 44.36 5.15 51.13 34.51
DEM (Ours) 56.94 28.24 54.34 7.78 53.31 40.22
Table 3: Summary results of the HELM evaluations on held-out scenarios, grouped by task-category for the 7B model. DEM outperforms data mixing approach in five out of six HELM task clusters.
Next, we combine the models fine tuned on single datasets with distribution vector and interpolation method using Eq 2 and 3 respectively. The corresponding results are shown in Table 2. We observe that both approaches perform significantly better than the best data mixing model for all 3 scenarios showing their effectiveness (see Appendix D for MMLU performance per domain). We also compare DEM with larger fine-tuned models (FlanPaLM (8B), OPT-IML (30B, 175B) and observe that DEM performs better although these models were trained on a larger mix of tasks and datasets compared to our model. Additionally, DEM - Distribution Vector performs better than DEM - Interpolation due to more flexible choice of $\omega_{i}$ (Section 4.2) and we use it in the rest of the paper.
We further expand our evaluation setting to include HELM scenarios. Here, we compare the performance of the pretrained model with DEM and data mixing model on multiple HELM held-out task clusters (see Table 3). DEM outperforms the data mixing approach in five out of six HELM task clusters. Surprisingly, for Math task category, the fine-tuned model performance degrades as compared to the pretrained model. Closer inspection reveals that this degradation is partly due to the fact that the instruction-tuned model does not output the answer in the correct format (as expected by HELM evaluation metric). The detailed HELM evaluation results (including results on ‘seen’ tasks) are reported in the Appendix D (Table 12).
6.2 Effect of Model Size Scaling
# Params Models MMLU BBH DROP MathQA
3B Data Mixing 41.08 31.36 25.98 28.54
DEM 43.67 34.14 28.89 29.78
7B Data Mixing 47.77 36.38 32.71 30.35
DEM 50.74 40.56 37.96 32.16
13B Data Mixing 52.7 40.48 43.15 30.72
DEM 54.53 42.65 46.59 33.13
Table 4: Effect of model size on the performance of the proposed approach. We observe performance improvement using DEM for both smaller (3B) and larger (13B) models compared to Data Mixing baseline.
We evaluate the performance of the proposed DEM approach with increasing model sizes using Open-LLaMA 3B, 7B, and 13B models, quantifying the impact with both smaller and larger models. We trained the baseline Data Mixing model using the method discussed in Section 5.5. On the other hand, we fine-tuned the models on each dataset separately and combined them using Eq 2, similar to the 7B model as discussed in Section 6.1. We use the same model mixing weight of $\omega_{i}=0.25$ (optimized for 7B model) for models of all sizes and present the results in Table 4. We observe that the model performance increases as we scale up the model size from 3B to 13B for both Data Mixing and DEM. Additionally, DEM yields performance improvement for each model size, showing the effectiveness and generalizability of the proposed approach with model size.
6.3 Impact of Different Training Datasets
In this section, we analyze the impact of each training dataset included in the mixture on the downstream task performance. For this, we progressively add the data distribution vector corresponding to each dataset to the base model (following Eq. 2) and evaluate the resulting model. We use $\omega_{i}=0.25$ for all datasets to keep the setup simple. The performance of the resulting models are presented in Table 5. We observe that these data sources yield different levels of performance gains, as expected. This can be due to the various levels of mismatch between the train and test distribution. We observe that combining the pretrained model with single-task distribution vectors (e.g Math QA) or smaller mix of tasks (e.g., CoT) leads to smaller improvement whereas large scale multi-task distribution vectors (e.g., P3 and SNI) yields a much larger performance gain, in comparison. It can also be due to the large diversity of tasks and samples in P3 and SNI. InstructDial is an exception, which can be due to the conversational nature of this dataset, making it very different from the evaluation tasks.
Training Dataset MMLU BBH DROP MathQA
Open LLaMA 40.31 32.84 24.38 27.71
+ CoT 41.30 33.68 25.46 28.44
+ MathQA 41.67 33.73 26.05 28.68
+ P3 47.12 36.58 30.82 30.35
+ InstDial 47.44 38.20 31.15 30.65
+ SNI 50.74 40.56 37.96 32.16
Table 5: Effect of progressively adding distribution vectors (Eq 1) from different data sources to the pretrained model using DEM (Eq 2). The performance increases as we add more data sources.
OpenLLaMA vs. Euclidean
P3 35.1
InstDial 85.1
SNI 34.1
CoT 3.2
MathQA 4.1
Data Mixing 74.6
DEM 20.8
Table 6: Euclidean distance between the base model (OpenLLaMA) and the fine-tuned models.
6.4 Properties of Distribution Vectors
To better understand the behavior of DEM, we examine the characteristics of the fine-tuned models and their corresponding distribution vectors, as defined in (Eq 1). We evaluate the similarity between models by calculating the Euclidean distance and the cosine similarity after converting their weights into a single flattened vector representation.
Individual model distance from base. In Table 6, we show the Euclidean distance from the base model to each fine-tuned model. Datasets with more examples (P3, Instruct Dial, and SNI) lead to models that are further away from the base. The largest change is caused by Instruct Dial (x3 compared to the second largest), since it introduces a very specific domain (i.e., conversations), and requires higher adaptation of the model. In contrast, smaller datasets (CoT, and Math QA) only contribute small changes (3-4 points). As expected, the distribution edited model (DEM) is closer to the base model than the models trained on the largest datasets. This is because DEM is derived from a weighted average of the individual vectors. Finally, we observe that the Data Mixing model has significantly higher euclidean distance (x3) from the base model compared to DEM, indicating that the Data Mixing approach introduces a larger change.
$\downarrow$ Dist. Vector $\rightarrow$ P3 InstDial SNI CoT MathQA
InstDial 0.07 -
SNI 0.09 0.08 -
CoT 0.02 0.01 0.02 -
MathQA 0.01 0.01 0.01 1.0 -
Data Mixing 0.27 0.29 0.19 0.10 0.10
DEM 0.44 0.87 0.43 0.06 0.06
Table 7: Cosine similarity between distribution vectors.
Figure 2: tSNE representation of the fine-tuning datasets. The centroids of the datasets are marked as larger points with captions.
Pairwise similarity between distribution vectors. Next, in Table 7 we compare the pairwise cosine similarity between the DVs from the fine-tuned models. We show that most of the individual DVs are almost orthogonal, except CoT and Math QA. This suggests that fine-tuning on these datasets does not lead to interference and introduces different abilities into the model.
To understand this, we sample 2,000 points from each dataset and plot their embedding representations into a common space using tSNE (van der Maaten and Hinton (2008), see Figure 2). [^2] We encode all texts after formatting them into their corresponding prompt using sentence-transformers/all-MiniLM-L12-v2. We observe a large number of CoT prompts that are close to the centroid of the MathQA dataset, which may explain the high similarity between their DVs. Note that CoT also has a small overlap with P3 but further away from their centroids, making the two DVs almost orthogonal. All other datasets have a minimal overlap between each other, and form independent clusters. We also study the relation between the combined DV and individual DVs (last row in Table 7). We observe that DEM is oriented towards the DVs from the models with the highest euclidean distance from the pretrained model.
Finally, we compare how the data mixing model is oriented relative to the individual DVs. The cosine similarity with all DVs is less than 0.3, however, the model is oriented towards the DVs of bigger and more diverse datasets (InstDail, P3, SNI). The similarity with the CoT and MathQA is a bit higher, but it still remains within 0.1.
Figure 3: Layer-wise Euclidean distance, comparison between the base OpenLLaMA model, and the tuned models. Darker colors mean higher absolute difference. The euclidean distance values are normalized per-model by the highest layer-distance for that model. The plots are invariant to the scale of the weight change.
Layer-wise distance of individual models from base. Finally, to fully understand the changes in the models and why DEM is an effective strategy for data distribution mixing, we zoom in even further into the layer-wise euclidean distance (Figure 3) between the individual task vectors and the base model (OpenLLaMA 7B). From Figure 3, it is evident that the changes in the tuned models occur mostly in the first three layers. The embedding layers remain relatively stable across different domains and dataset sizes, indicating that the fundamental properties are preserved. New knowledge is primarily acquired by the 2-3 layers, which contributes to the success of the proposed approach. Furthermore, this study suggests that when combining weights, it is not necessary to take into account all the weights involved. Instead, it is possible to safely remove or prune certain weights in the combination without significantly impacting the outcome, as also shown by Yadav et al. (2023).
6.5 Compute Cost Comparison
We use Eq. 4 to compare the real compute cost of the proposed DEM approach with the baseline data mixing method for 7B model on Nvidia A100 machines (with 8 gpus per node). Note that, this cost is specific to our setup and it can change depending on the model size, training parallelization scheme and other factors. In Table 8, we present the gpu-hours used by different training runs, as well as the validation runs needed for finding optimal model mixing weight $\omega_{i}$ in Eq 2. In each case, we did early stopping to obtain the best validation loss, which results in varying number of training steps (‘# steps’ in Table 8). As discussed in Appendix B, we use a combination of 10 weights to get the best model for DEM, which costs 23 gpu-hours. The total cost (training+validation) for DEM is 966 gpu-hours.
For the baseline data mixing, we trained 50 models with different weight combination (the exact weight selection process is described in Appendix B). Each run costs an average of 233 gpu-hours, resulting in a total cost of 11650 gpu-hours. This is more than 11 times the total cost of DEM.
Train/Val Runs time / step # steps # gpus Cost
DEM
- CoT 6.5 550 8 8
- Math QA 6.5 600 8 8.7
- P3 4.8 6000 32 256
- InstDial 5.2 23000 16 530
- SNI 5.24 6000 16 140
- Validation (10x) 2.1 500 8 23
Total 966
Data-Mixing (50x) 5.24 15000 16 11650
Table 8: Training cost (in gpu-hours) of 7B model on different instruction following datasets computed using Eq 4. Note that the number of steps is not equal to the number of examples.
7 Conclusions and Future Work
We proposed a simple and efficient approach for training on diverse data distributions that trains checkpoints individually on each data source and then combines them with basic element-wise vector operations. DEM siginficantly outperforms the standard weighted data mixing in terms of downstream performance and overall compute cost. Our experiments demonstrate that DEM works with both single-task (e.g. Math QA) and multi-task data distributions (e.g. SNI, P3), and that they can be incrementally added to the pretrained model, resulting in improved downstream performance. We further performed extensive analysis to better understand the properties of the learned distribution vectors, finding that DEM is better aligned with the individual models than baseline while remaining close to the original model.
In future, it is important to evaluate the proposed approach using other model architectures e.g. encoder-decoder or mixture of experts model to better understand its effectiveness with other model designs. Additionally, DEM can be further improved by using more sophisticated methods for combining the individual checkpoints that can reduce the negative effects of interfering data distributions.
Acknowledgments
We thank the anonymous reviewers for their helpful questions and comments, which have helped us improve the quality of the paper. We also want to thank Yang Li for their help in setting up HELM evaluation framework, and Thomas Müller and Lluís Màrquez for helpful discussions.
Limitations
While this paper proposes an efficient and effective alternative to data mixing for training multi-task and instruction-following models, it is important to acknowledge its limitations:
  • Task granularity. The distribution vectors of DEM are applicable to data distributions that span a single or multiple tasks. Our experimentation focused on existing data sources with different granularities ranging from several hundred tasks (e.g. P3) to a single one (e.g. MathQA), hence, the resulting distribution vectors captured varying task granularities. A detailed investigation of granularities and how to automatically group the data is an open area of investigation.
  • Architecture type. The proposed approach makes no specific assumptions regarding the architecture and should be, in principle, applicable to any architecture variant including Mixture-of-Expert models Fedus et al. (2022); Xue et al. (2024); Jiang et al. (2024); Sukhbaatar et al. (2024); Hu et al. (2024). Due to budget constraints, the evaluation of different architecture types was not included in the experiment plan. Therefore, the compatibility of DEM with different architecture types remains to be evaluated.
  • Storage Requirements. DEM reduces the computational cost of training models, but it requires storing a number of distribution vectors in the hard drive. For very large models, this creates the need for large storage capacity that may not be always available. One straight-forward solution to this problem is to use parameter-efficient methods to train the distribution vectors instead of full training or discard the distribution vectors once the optimal combination has been identified.
Appendix
Appendix A Training Hyperparameters
In this section, we describe the detailed hyperparameters that we used for fine tuning the OpenLLaMA model using different datasets separately for DEM and combined for Data Mixing and Concatenated Datasets. In all these cases, we use a constant learning rate of 2e-5 with a 2000 step warmup. We tested other learning rate schedules with cosine and linear decay in preliminary experiments, however, they lead to worse performance. We use AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay of 0.05 and gradient clipping of 1. We also adjust batch size for different datasets based on the validation loss. We use example packing to fit multiple training examples into a single sample of a batch for efficient training. This is a greedy packing approach where we pack training examples until we reach the max sequence length that we can fit into the model. We do not overflow an example into the next sample of a batch (as generally done during pretraining Brown et al. (2020)), rather use padding to fill the sample. The full setting is presented in Table 9. We show the batch size using total number of token after sample packing. The number of training steps indicates the step with best validation loss, and its variable for different datasets. Note that for InstDial, this value is particularly high because of the different kind of samples (i.e. dialog) that consists of this dataset.
Dataset / Batch Learning # Training
Method Size Rate Steps
CoT 65k 2e-5 550
MathQA 65k 2e-5 600
P3 262k 2e-5 6k
InstDial 131k 2e-5 23k
SNI 131k 2e-5 6k
Data Mixing 131k 2e-5 15k
Concatenated Datasets 131K 2e-5 17K
Table 9: Training hyperparameters for different models.
Appendix B Choosing Data Mixing Weights
Based on initial experiments, we determined the following hyperparameter ranges for the baseline data mixing approach – CoT: [0.05, 0.1, 0.15, 0.20], Math QA: [0.05, 0.1, 0.15, 0.20], P3: [0.25, 0.30, 0.35, 0.40], InstructDial: [0.30, 0.35, 0.40, 0.45], Super Natural Instructions: [0.15, 0.20, 0.25, 0.30]. Out of the 1024 possible weight combinations above, we randomly selected 50 combinations for training and selected the best weight setting based on validation-loss. The optimal data mixing setting was the following: P3 - 0.30, SNI - 0.20, Instruct-Dial - 0.40, MathQA - 0.05, CoT - 0.05 The total cost for this hyperparameter search procedure is listed in Table 8).
Appendix C Choosing DEM Weights
Models MMLU BBH DROP
Open LLaMA 40.31 32.84 24.38
DEM - Distribution Vector
$\omega=0.25$ 50.74 40.56 37.96
Random Search, x50 50.98 40.55 40.83
Table 10: Downstream task performance of the DEM w/ Distribution Vector. We compare the weight selection strategies: single-coefficient vs. random search with 50 iterations.
In order to select the optimal mixing weights for DEM - Distribution Vector (Eq 2), we perform a grid search over $\omega_{i}$ values. For each coefficient combination we evaluate the validation losses on the five datasets used for fine-tuning (Section 5.1), and select the model that minimizes their average. However, exhaustive grid search is expensive as the number of combinations grows exponentially. Thus, we simplify Eq 2 and optimize a single coefficient $\omega$ for all datasets. We found $\omega=0.25$ (out of 10 values) to produce the best validation loss and use it for all our experiments.
We chose the weights for DEM - Interpolation (Eq 3) in a similar manner as DEM - Distribution Vector by randomly sampling weights from the search grid and normalizing them to sum to 1. Additionally, we also tried the same weights as data mixing and equal weight of 0.2 for each of 5 datasets. The simplest strategy of equal weight performed on par with the best weight combination in terms of average val loss. So, we chose this and reported the corresponding results in Table 2
To measure the effect of using a single coefficient, we perform a limited budget experiment with 50 weight combinations, which are produced using individual weights for each distribution vector (Eq 1), sampled uniformly from the interval $[0,1]$. Our results show that the best single-coefficient models perform better or on par with the sampled models in terms of average validation loss. This formulation was also adopted in other model interpolation works Ilharco et al. (2022a); Yadav et al. (2023). In Table 10, we show the differences in performance on three benchmarks (MMLU, BBH, DROP) using the Open LLaMa 7B model. The two strategies have similar performance on MMLU and BBH but the random search has an advantage of 3 points on DROP. However, this increase comes at the expense of 5x increase in cost (10 evaluations for uniform vs. 50 evaluations for random search). The best distribution weights we found are: CoT - 0.1, InstDial - 0.12, MathQA - 0.1, P3 - 0.23, SNI - 0.45. We hypothesize that the single-vector weights will not be an optimal choice if there is high negative correlation between the vectors, i.e., the data distributions are conflicting.
Appendix D Fine-Grained Results
In Table 11 we show the model performance per domain on the MMLU benchmarking datasets. It covers five different categories, on all of which DEM outperforms the other alternatives.
In Table 12 we show the per-dataset results on HELM benchmark. We can see that our approach significantly outperforms data mixing and improves over the baseline model in most of the categories. Due to space limitations we show different datasets on different rows.
Training Dataset STEM Humanities Social Sciences Others Average
Open LLaMA v2 33.4 36.8 45.1 47.3 40.3
LLaMA 1 Touvron et al. (2023a) 34.0 30.5 38.3 38.1 35.1
LLaMA 2 Touvron et al. (2023b) 42.9 36.4 51.2 52.2 45.3
Public Pool of Prompts (P3) 25.4 32.9 44.2 41.2 35.7
Instruct Dial 31.9 37.8 44.5 43.5 39.3
Super Natural Instructions (SNI) 38.4 42.6 53.9 52.9 46.5
Chain of Thoughts (CoT) 34.4 38.3 47.1 48.1 41.7
Math QA 32.8 36.6 44.1 46.5 39.7
Data Mixing 39.2 44.8 55.6 52.6 47.8
Concatenated Datasets (1-5) 37.9 41.4 49.8 46.1 43.4
DEM - Interpolation (Ours) 39.7 47.2 58.5 56.2 50.1
DEM - Distribution Vector (Ours) 40.4 47.8 58.8 57.0 50.7
Table 11: MMLU domain specific task performance of models trained on different instruction following datasets (Section 5.1). We compare it with different pretrained and fine-tuned baselines (Section 5.5) and our proposed approach in Section 4.
Models MMLU BoolQ NarrativeQA NaturalQ closed NaturalQ open QUAC TruthfulQA IMDB CivilComments RAFT Wikifact
OpenLLaMA-v2 39.37 72.3 63.96 26.08 61.15 33.03 18.65 93.2 53.96 60.0 24.89
Data Mixing 43.96 85.3 71.24 21.84 19.5 34.52 42.35 87.0 64.8 69.09 21.94
DEM (ours) 46.61 82.4 71.24 28.1 69.39 40.22 29.82 96.6 53.75 67.27 26.82
Models ReasonAbstract ReasonNatural bABI Dyck GSM-8K Math-Eq Math-CoT LSAT Legal Imputation EntityMatch
OpenLLaMA-v2 18.51 21.1 45.25 52.0 5.5 12.58 8.33 20.43 48.67 81.66 83.89
Data Mixing 20.58 29.64 54.52 40.0 0.5 8.79 1.52 24.35 62.37 76.4 85.62
DEM (ours) 23.24 34.73 56.62 48.4 6.3 11.06 4.52 18.26 58.49 71.56 85.32
Table 12: Detailed HELM results on 22 scenarios. HELM datasets that are part of model-training (BoolQ, GSM-8K and IMDB), are excluded from the aggregated results presented in Table 3